Service Alerts Archive

Stay Informed

Subscribe to the OIT Service Status Mailing List
Receive an email message when an OIT service is down or service has degraded.

  • UPDATE 1011: Google Drive – Service Degradation, Reported on: January 17, 2017 8:52 am

    UPDATE 1011: Google is still working on the issue. Still no ETA. We are getting reports of general slowness and ‘connection reset’ errors when users try to access Google Drive. The issue is certainly upstream of us and is on Google’s end. Unfortunately our admins cannot resolve this issue, but rest… Read More »


    Status: Resolved
    Resolution:

    Service restored this afternoon.

    Updated on: January 17, 2017 6:23 pm


    Posted in: | Tags:
  • CBS (over-the-air channel 2) is presently down, Reported on: January 12, 2017 5:04 pm

    We are getting reports that CBS (over-the-air channel 2) is presently down. We’ve contacted our cable provider and they are aware of the issue. No ETA at this time.


    Status: Resolved
    Resolution:

    This issue was resolved.

    Updated on: January 13, 2017 10:40 am


    Posted in: | Tags:
  • Update 1715 – Webmail/ES mail inaccessible & users unable to send/receive in clients (i.e. Thunderbird), Reported on: January 6, 2017 4:48 pm

    At ~1620 today the OIT Help Desk began receiving reports of issues with accessing Webmail, and sending or receiving messages in email client software, like Thunderbird. The issue has been reported to OIT Email System Admins and they are working on resolving the issue now. No ETA as of this writing.


    Status: Resolved
    Resolution:

    As of 1715 all users of Webmail.uci.edu and ES Mail services should begin to see service restored.
    OIT Email system admins corrected and issue with a security certificate that was preventing users of these systems from connecting to their respective servers.

    Updated on: January 6, 2017 5:17 pm


    Posted in: | Tags:
  • Update 1435 – Power flux causing power outages in buildings across campus, Reported on: January 5, 2017 12:37 pm

    A power flux at 12:26 PM has caused various buildings in the Bio Sci and Physical Sciences to lose power. OIT Network Technicians are aware of the issue and are coordinating with Facilities, but do not yet have an ETA for restoration of services. UPDATE 1235: Buildings in the research park and some… Read More »


    Status: Resolved
    Resolution:

    OIT Network Technicians have reported that all network services have been restored. There may still be outstanding issues related to the power outage, but network issues have all been addressed. Please report any further issues to OIT Help Desk as normal.

    Updated on: January 5, 2017 2:55 pm


    Posted in: | Tags:
  • UPDATE 2218. EECS Servers – Service Degradation, Reported on: December 30, 2016 1:28 pm

    UPDATE 2218. Admins are working with the users to free up space on the server. UPDATE 0942. Our admins have found the source of the issue and are working with end users to resolve the issue. We are getting a couple of reports of low disk space on some of the EECS servers. Our administrators are awar… Read More »


    Status: Resolved
    Resolution:

    Admins are working with the users to free up space on the server.

    Updated on: January 2, 2017 10:21 pm


    Posted in: | Tags:
  • Service Degradation – FileNet/eDocs Scanning Service, Reported on: December 27, 2016 2:36 pm

    We are receiving reports that the the FileNet/eDocs Scanning Service is unavailable for some users. Individuals access this service via KFS (when trying to view attachments) and the SABE system (imaging system used by the UCIMC,) and these users reported seeing ‘invalid token’ errors. Our admins are… Read More »


    Status: Resolved
    Resolution:

    Service has been restored. We apologize for the trouble this may have caused.
    If you're still having issues, close and then re-open your browser. You may need to clear your cache and cookies as well.

    Updated on: December 27, 2016 2:50 pm


    Posted in: | Tags: